OpenStudy (anonymous):

FACTOR COMPLETELY: 100g^2h^2-x^2

OpenStudy (anonymous):

(10gh)^2 - (x)^2 use the identity

OpenStudy (anonymous):

Understood???

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

if u square 10gh u get 100g^2h^2 use identity a^2 - b^2 = (a+b) (a-b) n get yr answer

OpenStudy (anonymous):

a^2-b^2 = (a+b)(a-b) using this u get the answer as (10gh+x)(10gh-x)